Ask your own question, for FREE!
Mathematics 8 Online
zarkam21 (zarkam21):

Help :]

zarkam21 (zarkam21):

OpenStudy (anonymous):

for the probability orange candy = 1/5, while the probability for any other candy= 4/5 so the probability that exactly one of the candies is orange 1/5 * 4/5 = 4/25 = 0.16

OpenStudy (anonymous):

^ correct

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
1 day ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
1 day ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!